Please help me solve this question:
Solve using the substitution method:
r=-3s
r+4s=10
I don't understand how to write it out or what numbers to use.
Thanks,
Deb

Answer by Alan3354(69443)

Solve using the substitution method:
r=-3s
r+4s=10
--------
Sub for r in the 2nd equation
r + 4s = 10
-3s + 4s = 10
s = 10
--------
r = -3s
r = -30